25.00 mL aliquot of 0.016 M NaCl was titrated with 27.52 mL of 0.18 M

AgNO3. Determine whether a precipitate will form at end point. Ksp AgCl = 1.6

x10-10

 


1
Expert's answer
2021-08-16T02:45:59-0400

Solution:

Cinitial × Vinitial = Cfinal × Vfinal

Vfinal = V(NaCl) + V(AgNO3) = 25.00 mL + 27.52 mL = 52.52 mL

Vfinal = 52.52 mL


The concentration of NaCl when the solutions are mixed:

Cfinal(NaCl) = Cinitial(NaCl) × (Vinitial / Vfinal)

Cfinal(NaCl) = (0.016 M) × (25.00 mL / 52.52 mL) = 0.00762 M

Cfinal(NaCl) = 0.00762 M


The concentration of AgNO3 when the solutions are mixed:

Cfinal(AgNO3) = Cinitial(AgNO3) × (Vinitial / Vfinal)

Cfinal(AgNO3) = (0.18 M) × (27.52 mL / 52.52 mL) = 0.09432 M

Cfinal(AgNO3) = 0.09432 M


The concentration of Cl- when the solutions are mixed:

NaCl → Na+ + Cl-

According to the equation above:

[Cl-] = Cfinal(NaCl) = 0.00762 M


The concentration of Ag+ when the solutions are mixed:

AgNO3 → Ag+ + NO3-

According to the equation above:

[Ag+] = Cfinal(AgNO3) = 0.09432 M


AgCl(s) → Ag+(aq) + Cl-(aq)

The solubility product expression is as follows:

Ksp = [Ag+] × [Cl-] = 1.6×10-10

Ksp = 1.6×10-10

Calculate the ion product:

Q = [Ag+] × [Cl-] = (0.09432) × (0.00762) = 0.0007187 = 7.2×10-4

Q = 7.2×10-4


Compare Q with the Ksp.

If Q > Ksp, then AgCl will precipitate, but if Q < Ksp, it will not.

Because Q > Ksp, we predict that AgCl will precipitate when the two solutions are mixed.


Answer: Yes, AgCl will precipitate when the two solutions are mixed